OpenStudy (anonymous):

Determine whether -x^3/3x^2+5 is even, odd, or neither

jimthompson5910 (jim_thompson5910):

If you can show that f(-x) = f(x), then the function f(x) is even If you can show that f(-x) = -f(x), then the function f(x) is odd

OpenStudy (anonymous):

But I don't know how to solve.

jimthompson5910 (jim_thompson5910):

f(x) = (-x^3)/(3x^2+5) f(-x) = (-(-x)^3)/(3(-x)^2+5) f(-x) = (x^3)/(3x^2+5) Since f(-x) is NOT equal to f(x), we know that f(x) is NOT even

OpenStudy (anonymous):

It's odd?

jimthompson5910 (jim_thompson5910):

yep because f(x) = (-x^3)/(3x^2+5) -f(x) = -(-x^3)/(3x^2+5) -f(x) = (x^3)/(3x^2+5) which is equal to f(-x)

OpenStudy (anonymous):

Where did you get x^3?

jimthompson5910 (jim_thompson5910):

its in the original problem

OpenStudy (anonymous):

Sorry I was looking at something else/

OpenStudy (anonymous):

Can you explain your steps please?

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

I don't understand any of it.

jimthompson5910 (jim_thompson5910):

going from f(x) = (-x^3)/(3x^2+5) to f(-x) = (-(-x)^3)/(3(-x)^2+5) I replaced each 'x' with '-x', then simplified to get f(-x) = (x^3)/(3x^2+5)

OpenStudy (anonymous):

Okay I get it now. Thank you... and one other thing.

OpenStudy (anonymous):

How can I determine whether the equation y^2=6-x^2 defines y as a function of x

jimthompson5910 (jim_thompson5910):

solve for y, if you get one single equation for y, then the equation is a function if you get more than one equation, then the equation is not a function

OpenStudy (anonymous):

Solving for Y will I replace x with 0?

jimthompson5910 (jim_thompson5910):

no just isolate y

jimthompson5910 (jim_thompson5910):

in terms of x

jimthompson5910 (jim_thompson5910):

you want y = ...

OpenStudy (anonymous):

Right so at the moment it is y^2=6-x^2

jimthompson5910 (jim_thompson5910):

yep, then what

OpenStudy (anonymous):

I need to get rid of the square by using a square root?

jimthompson5910 (jim_thompson5910):

yep, that's correct, so what do you get when you do that

OpenStudy (anonymous):